# Model-Based Transfer Learning for Contextual Reinforcement Learning

### v1

@@ 4.1 Modeling assumptions
We make key assumptions about the task space, performance functions, and the generalization gap.
Assumption 1 (Continuity of the task space).
The task set $X$ is continuous, meaning that for any two tasks $x,x^{\prime}\in X$ and any $\epsilon>0$, there exists a $\delta>0$ such that $\|x-x^{\prime}\|<\delta$ implies the corresponding tasks $x$ and $x^{\prime}$ are indistinguishably close. Formally,
$\forall x,x^{\prime}\in X,\forall\epsilon>0,\exists\delta>0\text{ such that }\|x-x^{\prime}\|<\delta\implies\|J(x)-J(x^{\prime})\|<\epsilon.$
Assumption 2 (Smoothness of the performance function).
The performance function $J(x)$, $U(x)$, or $V(x)$ is smooth. This implies that these functions are continuously differentiable over $x$. Formally, $J(x),U(x),V(x)\in C^{1}(X)$, where $C^{1}(X)$ denotes the space of continuously differentiable functions on $X$.
@@
A linear model is used to model the generalization gap function, formally $\Delta\hat{J}(x,x^{\prime})\simeq\theta|x-x^{\prime}|$, where $\theta$ is the slope of the linear function and $x$ and $x^{\prime}$ are the context of the source task and target task, respectively.
Assumption 1 and 2 are for implementing MBTL, which relies on the continuous task space and smooth performance function. Moreover, we simplify the generalization gap function as a linear function inspired by empirical observations such as shown in Figure 1. Such structure (Assumption 3) enables the development of provable algorithms for solving SSTS. Algorithm 1 provides a generic psuedocode for a MBTL algorithm. The generalization gap structure is leveraged within Line 2 (function NextTask).
4.2 Naive strategies
To understand what is desirable in a method for solving SSTS, it is instructive to consider some naive strategies. An Equidistant Strategy (ES) samples source tasks so that the tasks are equidistant along the full range of tasks we aim to solve. In this method, source tasks are selected to be equidistant within the context space based on the transfer budget $K$. Mathematically, the selected source tasks are chosen such that the distances between consecutive tasks are approximately equal, i.e., $x_{k}=\frac{2k-1}{2K}|X|\quad\text{for }k=1,2,\ldots,K.$ Note that this strategy requires advanced knowledge of the total number of training tasks (e.g., training budget). A Greedy Strategy (GS) greedily selects the source task that is predicted to have the highest generalized performance in each round, assuming known and constant training performance. Figure 3 depicts the difference between ES and GS; observe that the normalized performance of the trained tasks is assumed to be 1. However, a critical question emerges concerning scenarios where training performance is unknown or stochastic. Furthermore, ES and GS are open-loop strategies; that is, they do not incorporate feedback from actual training outcomes.
@@ 4.3 Bayesian optimization
Bayesian optimization (BO) is a powerful strategy for finding the global optimum of an objective function when obtaining the function is costly or the function itself lacks a simple analytical form [26, 7]. BO integrates prior knowledge with observations to efficiently decide the next task to train, aiming to optimize the predicted value.
@@
Gaussian Process. Within the framework of BO, we model the source training performance $J_{\text{pred}}(x)$ using Gaussian Process (GP) regression. Specifically, the function $J_{\text{pred}}(x)$ is assumed to follow a GP $(J_{\text{pred}}(x)\sim\mathcal{GP}(m(x),k(x,x^{\prime})))$, where $m(x)$ represents the mean function, $\mathbb{E}[J_{\text{pred}}(x)]$, and $k(x,x^{\prime})$ is the covariance function, representing the expected product of deviations of $J_{\text{pred}}(x)$ and $J_{\text{pred}}(x^{\prime})$ from their respective means. Let $D_{k-1}$ denote the data observed up to iteration $k-1$, consisting of the pairs $\{(x_{i},V_{\text{obs}}(x_{i}))\}_{i=1,...,k-1}$, where $V_{\text{obs}}(x_{(1:i)})$ are the observed outcomes. The posterior prediction of $J_{\text{pred}}$ at a new point $x$, given the data $D_{k-1}$ and previous inputs $x_{(1:k-1)}$, is normally distributed as $P(J_{\text{pred},k}\mid D_{k-1})=\mathcal{N}(\mu_{k}(x),\sigma_{k}^{2}(x))$. $\mu_{k}(x)$ and $\sigma_{k}^{2}(x)$ are defined as $\mu_{k}(x)=m(x)+\mathbf{k}^{\top}(\mathbf{K}+\sigma^{2}\mathbf{I})^{-1}\mathbf{y}$ and $\sigma_{k}^{2}(x)=k(x,x)-\mathbf{k}^{\top}(\mathbf{K}+\sigma^{2}\mathbf{I})^{-1}\mathbf{k}$, with $\mathbf{k}$ being the vector of covariances between $x$ and each $x_{i}$ in the observed data, i.e., $\mathbf{k}=[k(x,x_{1}),\ldots,k(x,x_{k-1})]$, and $\mathbf{K}$ is the covariance matrix for the observed inputs, defined as $\mathbf{K}=[k(x_{i},x_{j})]_{1\leq i,j\leq k-1}$. This formulation enables the GP to update its beliefs about the posterior prediction with every new observation, progressively improving the estimation.
@@
Acquisition function. The acquisition function plays a critical role in BO by guiding the selection of the next source training task. At each decision step $k$, the task $x_{k}$ is chosen by maximizing the acquisition function, as denoted by $x_{k}=\operatorname*{arg\,max}_{x}a(x;x_{(1:k-1)})$. Especially in our case, the acquisition function can be designed as the expected marginal improvement of generalized performance across all tasks. One effective strategy employed in the acquisition function is the upper confidence bound (UCB) acquisition function, which considers the trade-off between the expected performance of a task based on current models (exploitation) and the measure of uncertainty associated with the task’s outcome (exploration). It is defined as follows:
$a(x;x_{(1:k-1)})=\mathbb{E}_{x^{\prime}\in X}[[\mu_{k-1}(x)+\beta_{k}^{1/2}\sigma_{k-1}(x)-\Delta J(x,x^{\prime})-U(x^{\prime};x_{(1:k-1)})]_{+}]$ (4)
@@
where $[\cdot]_{+}$ represents $\max(\cdot,0)$ and we can use various forms of $\beta_{k}$.
Generalized performance. We focus on the generalized performance evaluated across all target tasks, defined as $g(x)=\mathbb{E}_{x^{\prime}\in X}[J_{\text{pred}}(x)-\Delta J(x,x^{\prime})]=J_{\text{pred}}(x)-\mathbb{E}_{x^{\prime}\in X}[\Delta J(x,x^{\prime})]$.
@@ 4.4 Regret analysis
We use regret to quantify the effectiveness of our source task selection based on BO. Specifically, we define regret at iteration $k$ as $r_{k}=g(x^{*}_{k})-g(x_{k})$, where $g(x^{*}_{k})$ represents the maximum generalized performance achievable across all tasks, and $g(x_{k})$ is the generalized performance at the current task selection $x_{k}$. Consequently, the cumulative regret after $K$ iterations is given by $R_{K}=\sum_{k=1}^{K}r_{k}$, summing the individual regrets over all iterations. Following the framework presented by Srinivas et al. [35], our goal is to establish that this cumulative regret grows sublinearly with respect to the number of iterations. Mathematically, we aim to prove that $\lim_{K\rightarrow\infty}\frac{R_{K}}{K}=0$, indicating that, on average, the performance of our strategy approaches the optimal performance as the number of iterations increases.
Regret of MBTL-GP. Having established the general framework for regret analysis, we now turn our attention to the specific regret properties of our MBTL-GP algorithm. To analyze the regret of MBTL-GP, consider the scaling factor for the UCB acquisition function given by $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$ in Equation 4. It is designed to achieve sublinear regret with high probability, aligning with the theoretical guarantees outlined in Theorem 1 and 5 from [35].
@@
Impact of search space elimination. In this section, we demonstrate that strategic reduction of the possible sets, guided by insights from previous task selections or source task training performance, leads to significantly tighter regret bounds than Theorem 1. By focusing on the most promising regions of the task space, our approach enhances learning efficiency and maximizes the model’s performance and applicability. Given the generalization gap observed in Figure 1, we observe that performance loss decreases as the context similarity increases. We model the degradation from the source task using a simple linear function in Assumption 3. Training on the source task can solve a significant portion of the remaining tasks. Our method progressively eliminates partitions of the task space at a certain rate with each iteration. If the source task selected in the previous steps could solve the remaining target task sufficiently, we can eliminate the search space at a desirable rate. Consequently, at each step, we effectively reduce the search space.
Given the generalization gap observed in Figure 1, we model the degradation from the source task using a simple linear function in Assumption 3. While the figure might not strictly appear linear, the linear approximation simplifies analysis and is supported by empirical observations. Training on the source task can solve a significant portion of the remaining tasks. Our method progressively eliminates partitions of the task space at a certain rate with each iteration. If the source task selection in the previous step sufficiently addresses the remaining target tasks, we can reduce the search space at a desirable rate. Consequently, at each step, we effectively focus on a reduced search space.
Definition 2 (Reduced search space).
We define the reduced search space $X_{k}$ as a subset of $X$, with each element $x^{\prime}\in X_{k}$, such that $U(x^{\prime};x_{(1:k-1)})\leq\hat{J}(x_{k})+\Delta J(x^{\prime},x_{k})$.
@@
Figure 4: Empirical results of the restriction of search space by MBTL-GP compared to two examples from Corollaries 2.1 and 2.2.
@@
Proofs for Corollaries 2.1 and 2.2 are provided in Appendix A.1.3 and A.1.4, respectively. Based on our experiments presented in Section 5, the rate of elimination of the largest segment for MBTL-GP is shown in Figure 4.

### v2

@@ 4.1 Modeling assumptions
We consider a task set $X$ that is continuous and the performance function $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi},x),V(x)$ for a policy $\pi$ to be smooth over the task space $X$. In practice, such as control systems, tasks often vary continuously and smoothly rather than abruptly. For example, adjusting the angle of a robotic arm by a small amount typically results in a small change in the system and optimal action.
Inspired by the empirical generalization gap performance as observed in Figure 1, we estimate the generalization gap $\Delta\hat{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x^{\prime})$ with a linear function.
@@
A linear function is used to model the generalization gap function, formally $\Delta\hat{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x^{\prime})\simeq\theta|x-x^{\prime}|$, where $\theta$ is the slope of the linear function and $x$ and $x^{\prime}$ are the context of the source task and target task, respectively.
The relaxation of this assumption can yield additional efficiency benefits but also adds modeling complexity, and thus is an interesting direction for future work.
@@ 4.2 Bayesian optimization
Bayesian optimization (BO) is a powerful strategy for finding the global optimum of an objective function when obtaining the function is costly, or the function itself lacks a simple analytical form [31, 7]. BO integrates prior knowledge with observations to efficiently decide the next task to train by using the acquisition function. MBTL is a BO method which optimizes for promising source tasks by leveraging Assumption 1 in its acquisition function. The role of BO is to approximate ${V}(x_{k};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$ (see Equation 3) using the data acquired thus far. The next source task $x_{k}$ is then selected using this estimate.
@@
Gaussian process (GP) regression. Within the framework of BO, we model the source training performance ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ \forall x\in X\setminus x_{1:k}}$ using Gaussian process (GP) regression. Specifically, the function ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)$ is assumed to follow a GP $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)\sim\mathcal{GP}(\mathbb{E}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)],k(x,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})))$, where $k(x,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})$ is the covariance function, representing the expected product of deviations of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{\tilde{x}},}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})$ from their respective means. Let $D_{k-1}$ denote the data observed up to iteration $k-1$, consisting of the pairs $\{(x_{i},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J(\pi_{i},x_{i})})\}_{i=1,...,k-1}$. The estimated performance $\hat{J}_{k}$ after querying $k-1$ samples is updated as more samples are obtained. The posterior prediction of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}_{k}}$ at a new point $x$, given the data $D_{k-1}$ and previous inputs ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1:k-1}}$, is normally distributed as $P(\hat{J}_{k}\mid D_{k-1})=\mathcal{N}(\mu_{k}(x),\sigma_{k}^{2}(x))$. $\mu_{k}(x)$ and $\sigma_{k}^{2}(x)$ are defined as $\mu_{k}(x)=\mathbb{E}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{J}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)]+\mathbf{k}^{\top}(\mathbf{K}+\sigma^{2}\mathbf{I})^{-1}\mathbf{y}$ and $\sigma_{k}^{2}(x)=k(x,x)-\mathbf{k}^{\top}(\mathbf{K}+\sigma^{2}\mathbf{I})^{-1}\mathbf{k}$, with $\mathbf{k}$ being the vector of covariances between $x$ and each $x_{i}$ in the observed data, i.e., $\mathbf{k}=[k(x,x_{1}),\ldots,k(x,x_{k-1})]$, and $\mathbf{K}$ is the covariance matrix for the observed inputs, defined as $\mathbf{K}=[k(x_{i},x_{j})]_{1\leq i,j\leq k-1}$. This enables the GP to update its beliefs about the posterior prediction with every new observation, progressively improving the estimation.
@@
Acquisition function. The acquisition function plays a critical role in BO by guiding the selection of the next source training task. At each decision step $k$, the task $x_{k}$ is chosen by maximizing the acquisition function, as denoted by $x_{k}=\operatorname*{arg\,max}_{x}a(x;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1:k-1}})$. One effective strategy employed in the acquisition function is the upper confidence bound (UCB) acquisition function, which considers the trade-off between the expected performance of a task based on the current task (exploitation) and the measure of uncertainty associated with the task’s outcome (exploration) [40]. Especially in our case, the acquisition function can be designed as UCB function subtracted by generalization gap and so-far best performance. It is defined as follows:
$a(x;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1:k-1}})=\mathbb{E}_{x^{\prime}\in X}[[\mu_{k-1}(x)+\beta_{k}^{1/2}\sigma_{k-1}(x)-\Delta\hat{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})-{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{{1:k-1}},x^{\prime}})]_{+}]$ (4)
@@
where $[\cdot]_{+}$ represents $\max(\cdot,0)$ and we can use various forms of $\beta_{k}$, which is the trade-off parameter between exploitation and exploration.
@@ 4.3 Regret analysis
We use regret to quantify the effectiveness of our source task selection based on BO. Specifically, we define regret at iteration $k$ as $r_{k}=\hat{V}(x_{k}^{*};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})-\hat{V}(x_{k};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$, where $\hat{V}(x_{k}^{*};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$ represents the maximum generalization performance achievable across all tasks, and $\hat{V}(x_{k};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$ is the performance at the current task selection $x_{k}$. Consequently, the cumulative regret after $K$ iterations is given by $R_{K}=\sum_{k=1}^{K}r_{k}$, summing the individual regrets over all iterations. Following the framework presented by Srinivas et al. [40], our goal is to establish that this cumulative regret grows sublinearly with respect to the number of iterations. Mathematically, we aim to prove that $\lim_{K\rightarrow\infty}\frac{R_{K}}{K}=0$, indicating that, on average, the performance of our strategy approaches the optimal performance as the number of iterations increases.
Regret of MBTL. Having established the general framework for regret analysis, we now turn our attention to the specific regret properties of our MBTL algorithm. To analyze the regret of MBTL, consider the scaling factor for the UCB acquisition function given by $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$ in Equation 4. It is designed to achieve sublinear regret with high probability, aligning with the theoretical guarantees outlined in Theorem 1 and 5 from [40].
@@
Impact of search space elimination. In this section, we demonstrate that strategic reduction of the possible sets, guided by insights from previous task selections or source task training performance, leads to significantly tighter regret bounds than Theorem 1. By focusing on the most promising regions of the task space, our approach enhances learning efficiency and maximizes the policy’s performance and applicability. Given the generalization gap observed in Figure 1, we observe that performance loss decreases as the context similarity increases. We model the degradation from the source task using a linear function in Assumption 1. Training on the source task can solve a significant portion of the remaining tasks. Our method progressively eliminates partitions of the task space at a certain rate with each iteration. If the source task selected in the previous steps could solve the remaining target task sufficiently, we can eliminate the search space at a desirable rate. Formally, we can define the search space at $k$-th iteration as follows:
Definition 2 (Search space).
We define the search space $X_{k}$ at iteration $k$ as a subset of $X$, with each element $x^{\prime}\in X_{k}$, such that $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1},x^{\prime}})\leq\hat{J}(\pi_{x_{k}},x^{\prime}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\Delta J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x_{k}},x^{\prime}})$.
Given the generalization gap observed in Figure 1, we model the degradation from the source task using a linear function in Assumption 1. While the figure might not strictly appear linear, the linear approximation simplifies analysis and is supported by empirical observations. Training on the source task can solve a significant portion of the remaining tasks. Our method progressively eliminates partitions of the task space at a certain rate with each iteration. If the source task selection in the previous step sufficiently addresses the remaining target tasks, we can reduce the search space at a desirable rate. Consequently, at each step, we effectively focus on a reduced search space.
@@
Figure 3: Empirical results of the restriction of search space by MBTL compared to two examples from Corollaries 2.1 and 2.2.
@@
Proofs for Corollaries 2.1 and 2.2 are provided in Appendix A.3.2 and A.3.3, respectively. Based on our experiments presented in Section 5, the rate of elimination of the largest segment for MBTL is shown in Figure 3.